OpenStudy (sleepyjess):

Write the expression in standard form. \(\large\dfrac{6}{4-13i}\)

OpenStudy (sleepyjess):

\(\large -\dfrac{24}{185}~-~\dfrac{78}{185}i\) \(\large \dfrac{24}{185}~+~\dfrac{78}{185}i\) \(\large -\dfrac{24}{185}~+~\dfrac{78}{185}i\) \(\large \dfrac{24}{185}~-~\dfrac{78}{185}i\)

OpenStudy (sleepyjess):

@thomaster @ganeshie8 @satellite73

OpenStudy (anonymous):

\[\large\dfrac{6}{4-13i}\times \frac{4+13i}{4+13i}\] is a start

OpenStudy (anonymous):

the denominator will be \(4^2+13^2=185\)

OpenStudy (anonymous):

which is the whole point of multiplying by the conjugate the numerator is whatever you get

OpenStudy (sleepyjess):

why \(13^2\) wouldn't it be \(13i^2\)

OpenStudy (anonymous):

\[(a+bi)(a-bi)=a^2+abi-abi-(bi)^2=a^2+b^2\]

OpenStudy (anonymous):

in brief \((a+bi)(a-bi)=a^2+b^2\) a real number

OpenStudy (sleepyjess):

I got \(\large \dfrac{-54}{185}\)

OpenStudy (sleepyjess):

so would it be d?